OpenStudy (anonymous):

Suppose 60% of kids who visit a doctor have a fever, and 30% of kids with a fever have sore throats. What's the probability that a kid who goes to the doctor has a fever and a sore throat?

OpenStudy (turingtest):

To find compound probabilities like this multiply the individual probabilities. For instance toss a coin twice what are the odds of heads both times? (1/2)(1/2)=1/4 So rewrite your percentages as fractions or decimals and multiply to get your answer.

OpenStudy (anonymous):

So the answer would be 0.18?

OpenStudy (turingtest):

yes, though they probably want it as a percentage :)

OpenStudy (anonymous):

Okie doke, thank you :)

OpenStudy (anonymous):

18% if you dont know for sure the answer, use a number to confirm. in this case, say 100 kids go to the doctor, 60 has fever, 30% of the fever kids has sore throat. 30% of 60 is 18. simple

OpenStudy (anonymous):

(6/10) * (3/10) = 18/100 = 9/50 = 18%